OpenStudy (anonymous):

Integral from 0 to 1/2 x sin^-1 (x) dx

OpenStudy (anonymous):

@sirm3d

OpenStudy (anonymous):

Integration by parts, and some subs, switching boundaries.trigs..all in 1 :D

OpenStudy (anonymous):

and brb, washroom

OpenStudy (sirm3d):

\[\int\limits_{0}^{1}x \sin^{-1} x dx\] the choice is \[u=sin^{-1}x, \quad dv = x dx\]\[du=\frac{1}{\sqrt {1-x^2}},\;v=\frac{1}{2}x^2\]

OpenStudy (sirm3d):

type error. upper limit of integration=1/2

OpenStudy (anonymous):

My work, but dont think u can understand it :P

OpenStudy (anonymous):

I was stucked in x^2 / 2 sqrt (1-x^2) for at least 15mins.. hard to decide to use subs, parts, trig, or whatever..

OpenStudy (sirm3d):

seeing \[\sqrt{a^2-x^2}\quad \sqrt{a^2+x^2}\quad \sqrt{x^2-a^2}\] calls for trigonometric substitution EVEN IF a simple substitution will work, because a trigonometric substitution always work.

OpenStudy (anonymous):

There is one w/ sqrt (1 - w^2) < this is much harder to use trig subs

OpenStudy (anonymous):

The problem is, I only have 6 mins per ques on the exam >.>

OpenStudy (sirm3d):

for the limits, \(x=\sin \theta\) \[x=0: \theta = 0\]\[x=1/2: \theta = \pi/6\]

OpenStudy (anonymous):

@lovekblue u have dne the right substitution after substitution u will get integral of 1/2 sin^2θ dθ

OpenStudy (sirm3d):